kenzie126:

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 7 over 8 power, all over 2 to the 1 over 4 power A- the eighth root of 2 to the fifth power B- the fifth root of 2 to the eighth power C- the square root of 2 to the 5 over 8 power D- the fourth root of 2 to the sixth power

jhonyy9:

use this \[x^{1/3} = \sqrt[3]{x}\]

jhonyy9:

\[\frac{ 2^{7/8} }{ 2^{1/4} } = ?\]

jhonyy9:

and use this \[\frac{ a^{x} }{ a^{y} } = a^{(x-y)}\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[\frac{ 2^{7/8} }{ 2^{1/4} } = ?\] \(\color{#0cbb34}{\text{End of Quote}}\) how can you re-write this using the above written formula ?

jhonyy9:

a = 2 x = 7/8 y = 1/4

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 and use this \[\frac{ a^{x} }{ a^{y} } = a^{(x-y)}\] \(\color{#0cbb34}{\text{End of Quote}}\) substitute them inside this formula

jhonyy9:

@kenzie126 do you understand ?

jhonyy9:

@supie any idea ?